# Language Models Represent Space and Time

### v1

@@ Abstract
The capabilities of large language models (LLMs) have sparked debate over whether such systems just learn an enormous collection of superficial statistics or a coherent model of the data generating process—a world model. We find evidence for the latter by analyzing the learned representations of three spatial datasets (world, US, NYC places) and three temporal datasets (historical figures, artworks, news headlines) in the Llama-2 family of models. We discover that LLMs learn linear representations of space and time across multiple scales. These representations are robust to prompting variations and unified across different entity types (e.g. cities and landmarks). In addition, we identify individual “space neurons” and “time neurons” that reliably encode spatial and temporal coordinates. Our analysis demonstrates that modern LLMs acquire structured knowledge about fundamental dimensions such as space and time, supporting the view that they learn not merely superficial statistics, but literal world models.
@@ 1 Introduction
In this work, we take the question of whether LLMs form world (and temporal) models as literally as possible—we attempt to extract an actual map of the world! Specifically, we construct six datasets containing the names of places or events with corresponding space or time coordinates that span multiple spatiotemporal scales: locations within the whole world, the United States, and New York City in addition to the death year of historical figures from the past 3000 years, the release date of art and entertainment from 1950s onward, and the publication date of news headlines from 2010 to 2020. Using the Llama-2 family of models (Touvron et al., 2023), we train linear regression probes (Alain & Bengio, 2016; Belinkov, 2022) on the internal activations of the names of these places and events at each layer to predict their real-world location or time.
These probing experiments reveal evidence that models build spatial and temporal representations throughout the early layers before plateauing at around the model halfway point with larger models consistently outperforming smaller ones (§ 3.1). We then show these representations are (1) linear, given that nonlinear probes do not perform better (§ 3.2), (2) fairly robust to changes in prompting (§ 3.3), and (3) unified across different kinds of entities (e.g. cities and natural landmarks).
One possible explanation for our results is that the model only learns a mapping from place to a country (for example) and it is the probe that actually learns the global geometry of how these different groups are geospatially (or temporally) related. To study this, we conduct a series of robustness checks to understand how the probes generalize across different data distributions (§ 4.1) and how probes trained on the PCA components perform (§ 4.2). Our findings suggest that the probes memorize the absolute positioning of these groups, but that the model does have representations which reflect the relative positioning. In other words, the probe learns the mapping from model coordinates to human interpretable coordinates. Finally, we use our probes to find individual neurons which activate as a function of space or time, providing strong evidence that the model is truly using these features (§ 5).
@@ 2.3 Evaluation
To evaluate the performance of our probes we report standard regression metrics such as $R^{2}$ and Spearman rank correlation on our test data (correlations averaged over latitude and longitude for spatial features). An additional metric we compute is the proximity error for each prediction, defined as the fraction of predictions closer to the target point than the actual prediction. The intuition is that for spatial data, absolute error metrics can be misleading (a 500km error for a city on the East Coast of the United States is far more significant than a 500km error in Siberia), so when analyzing errors per prediction, we often report this metric to account for the local differences in desired precision.
@@ 3.1 Existence
We first investigate the following empirical questions: do models represent time and space at all? If so, where internally in the model? Does the representation quality change substantially with model scale? In our first experiment, we train probes for every layer of Llama-2-{7B, 13B, 70B} for each of our space and time datasets. Our main results, depicted in Figure 2, show fairly consistent patterns across datasets. In particular, both spatial and temporal features can be recovered with a linear probe, these representations are more accurate with increasing model scale, and the representations smoothly increase in quality throughout the first half of the layers of the model before reaching a plateau. These observations are consistent with results from the factual recall literature demonstrating that early-to-mid MLP layers are responsible for recalling information about factual subjects (Meng et al., 2022a; Geva et al., 2023).
The dataset with the worst performance is the New York City dataset. This was expected given the relative obscurity of most of the entities compared with other datasets. However, this is also the dataset where the largest model has the best relative performance, at nearly 2x the $R^{2}$ of smaller models, suggesting that sufficiently large LLMs could eventually form detailed spatial models of individual cities.
@@ 3.2 Linear Representations
Within the interpretability literature, there is a growing body of evidence supporting the linear representation hypothesis that features within neural networks are represented linearly, that is, the presence or strength of a feature can be read out by projecting the relevant activation on to some feature vector (Mikolov et al., 2013; Olah et al., 2020; Elhage et al., 2022b). However, these results are almost always for binary or categorical features, unlike the naturally continuous features of space or time.
@@ 4.1 Verification via Generalization
4.1.1 Block holdout generalization
@@
We find that while generalization performance suffers, especially for the spatial datasets, it is clearly better than random. By plotting the predictions of the held out states or countries in Figures 10 and 11, a qualitatively clearer picture emerges. That is, the probe correctly generalizes by placing the points in the correct relative position (as measured by the angle between the true and predicted centroid) but not in their absolute position. We take this as weak evidence that the probes are extracting explicitly learned features by the model, but are memorizing the transformation from model coordinates to human coordinates. However, this does not fully rule out the underlying binary features hypothesis, as there could be a hierarchy of such features that do not follow country or decade boundaries.
@@
4.1.2 Cross entity generalization
@@
Similar to the above, we distinguish these hypotheses by training a series of probes where the train-test split is performed to hold out all points of a particular entity class  [^1] We only do this entities which do not make up the majority of the training data (e.g., as is the case with populated places for the world dataset and songs for the entertainment dataset) which is partially responsible for the discrepancies in the nominal cases for Tables 3 and 4. Table 4 reports the proximity error for the entities in the default test split compared to when heldout, averaged over all such splits as before. The results suggest that the probes largely generalize across entity types, with the main exception of the entertainment dataset [^2] We note in this case the Spearman correlation is still high, suggesting this is an issue with bias generalization, as the different entity types are not uniformly distributed in time..
@@ 6 Related Work
Our work is most directly inspired by prior research into the extent to which deep learning systems form interpretable models of their data generating process. The cleanest demonstrations have from come from GPT style models trained on chess (Toshniwal et al., 2022) and Othello games (Li et al., 2022) which were shown to have explicit representations of the board and game state. Subsequent work on Othello has shown these representations are also linear (Nanda et al., 2023). In true language models, Li et al. (2021) show that an entity’s dynamic properties or relations can be linearly read out from representations at different points in the context. Abdou et al. (2021) and Patel & Pavlick (2021) show LLMs have representations that reflect perceptual and conceptual structure in color and spatial (but not geographic) domains, analogous to our attempts to show representations reflect geographic structure. Most related to our work is Konkol et al. (2017) and Liétard et al. (2021) who study the degree to which geography is reflected in word embeddings or small language models. Liétard et al. (2021) conclude the amount of information is “limited” but that larger models show more signs of geographic knowledge. By studying models with 100x the parameters on 10x the labeled data, we confirm this hypothesis.
@@
Intepretability
More broadly, our work draws upon many results and ideas from the interpretability literature (Räuker et al., 2023), especially in topics related to probing (Belinkov, 2022), BERTology (Rogers et al., 2021), the linearity hypothesis and superposition (Elhage et al., 2022b), and mechanistic interpretability (Olah et al., 2020). More specific results related to our work include Hanna et al. (2023) who find a circuit implementing greater-than in the context of years, and Goh et al. (2021) who find neurons in multimodal models corresponding to places, similar to our time and space neurons.
@@ 7 Discussion
We have provided evidence that LLMs learn linear representations of space and time that are unified across entity types and fairly robust to prompting, and that there exists individual neurons that are highly sensitive to these features. The corollary is that next token prediction alone is sufficient for learning a literal map of the world give sufficient model and data size.
Our analysis raises many interesting questions for future work. While we showed that it is possible to linearly reconstruct a sample’s absolute position in space or time, and that some neurons use these probe directions, the true extent and structure of spatial and temporal representations remain unclear. In particular, we conjecture that the most canonical form of this structure is a discretized hierarchical mesh, where any sample is represented as a linear combination of its nearest basis points at each level of granularity. Moreover, the model can and does use this coordinate system to represent absolute position using the correct linear combination of basis directions in the same way a linear probe would. We expect that as models scale, this mesh is enhanced with more basis points, more scales of granularity (e.g. neighborhoods in cities), and more accurate mapping of entities to model coordinates (Michaud et al., 2023).
This suggests future work on extracting representations in the model’s coordinate system rather than trying to reconstruct human interpretable coordinates, perhaps with sparse autoencoders (Cunningham et al., 2023). Another confounder in our analysis, and factual recall research more broadly, is the existence of many entities in our dataset which the model is unaware of, contaminating our activation datasets. We would be interested in methods that can identify when a model recognizes a particular entity beyond simply prompting for specific facts and risking hallucinations.
We also barely scratched the surface of understanding how these spatial and temporal world models are learned, recalled, or used internally. By looking across training checkpoints, it may be possible to localize a point in training when a model organizes constituent is_in_place_X features into a coherent geometry or else conclude this process is gradual (Liu et al., 2021). We expect that the model components which construct these representations are similar or identical to those for factual recall (Meng et al., 2022a; Geva et al., 2023). In preliminary experiments, we found our models had trouble answering basic spatial and temporal relations questions without relying on multi-step reasoning, complicating any causal intervention analysis Wang et al. (2022), but think that this is the natural next step for understanding when and how these features are used.
Finally, we note that the representation of space and time has received much more attention in biological neural networks than artificial ones (Buzsáki & Llinás, 2017). Place and grid cells (O’Keefe & Dostrovsky, 1971; Hafting et al., 2005) in particular are among the most well-studied in the brain, as we expect it to be productive source of inspiration for future work on LLMs.
@@ Appendix A Datasets
Appendix B Additional Results

### v3

@@ Abstract
The capabilities of large language models (LLMs) have sparked debate over whether such systems just learn an enormous collection of superficial statistics or a set of more coherent and grounded representations that reflect the real world. We find evidence for the latter by analyzing the learned representations of three spatial datasets (world, US, NYC places) and three temporal datasets (historical figures, artworks, news headlines) in the Llama-2 family of models. We discover that LLMs learn linear representations of space and time across multiple scales. These representations are robust to prompting variations and unified across different entity types (e.g. cities and landmarks). In addition, we identify individual “space neurons” and “time neurons” that reliably encode spatial and temporal coordinates. While further investigation is needed, our results suggest modern LLMs learn rich spatiotemporal representations of the real world and possess basic ingredients of a world model.
@@ 1 Introduction
In this work, we take the question of whether LLMs form world (and temporal) models as literally as possible—we attempt to extract an actual map of the world! While such spatiotemporal representations do not constitute a dynamic causal world model in their own right, having coherent multi-scale representations of space and time are basic ingredients required in a more comprehensive model.
Specifically, we construct six datasets containing the names of places or events with corresponding space or time coordinates that span multiple spatiotemporal scales: locations within the whole world, the United States, and New York City in addition to the death year of historical figures from the past 3000 years, the release date of art and entertainment from 1950s onward, and the publication date of news headlines from 2010 to 2020. Using the Llama-2 (Touvron et al., 2023) and Pythia Biderman et al. (2023) family of models, we train linear regression probes (Alain & Bengio, 2016; Belinkov, 2022) on the internal activations of the names of these places and events at each layer to predict their real-world location (i.e., latitude/longitude) or time (numeric timestamp).
These probing experiments reveal evidence that models build spatial and temporal representations throughout the early layers before plateauing at around the model halfway point with larger models consistently outperforming smaller ones (§ 3.1). We then show these representations are (1) linear, given that nonlinear probes do not perform better (§ 3.2), (2) fairly robust to changes in prompting (§ 3.3), and (3) unified across different kinds of entities (e.g. cities and natural landmarks). We then conduct a series of robustness checks to understand how our probes generalize across different data distributions (§ 4.1) and how probes trained on the PCA components perform (§ 4.2). Finally, we use our probes to find individual neurons which activate as a function of space or time and use basic causal interventions to verify their importance in spatiotemporal modeling, providing strong evidence that the model is truly using these features (§ 5).
@@ 2.3 Evaluation
To evaluate the performance of our probes we report standard regression metrics such as $R^{2}$ and Spearman rank correlation on our test data (correlations averaged over latitude and longitude for spatial features). An additional metric we compute is the proximity error for each prediction, defined as the fraction of entities predicted to be closer to the target point than the prediction of the target entity. The intuition is that for spatial data, absolute error metrics can be misleading (a 500km error for a city on the East Coast of the United States is far more significant than a 500km error in Siberia), so when analyzing errors per prediction, we often report this metric to account for the local differences in desired precision.
@@ 3.1 Existence
We first investigate the following empirical questions: do models represent time and space at all? If so, where internally in the model? Does the representation quality change substantially with model scale? In our first experiment, we train probes for every layer of Llama-2-{7B, 13B, 70B} and Pythia-{160M, 410M, 1B, 1.4B, 2.8B, 6.9B} for each of our space and time datasets. Our main results, depicted in Figure 2, show fairly consistent patterns across datasets. In particular, both spatial and temporal features can be recovered with a linear probe, these representations smoothly increase in quality throughout the first half of the layers of the model before reaching a plateau, and the representations are more accurate with increasing model scale. The gap between the Llama and Pythia models is especially striking, and we suspect is due to the large difference in pre-training corpus size (2T and 300B tokens respectively). For this reason, we report the rest of our results on just the Llama models.
The dataset with the worst performance is the New York City dataset. This was expected given the relative obscurity of most of the entities compared with other datasets. However, this is also the dataset where the largest model has the best relative performance, suggesting that sufficiently large LLMs could eventually form detailed spatial models of individual cities.
@@ 3.2 Linear Representations
Within the interpretability literature, there is a growing body of evidence supporting the linear representation hypothesis that features within neural networks are represented linearly, that is, the presence or strength of a feature can be read out by projecting the relevant activation on to some feature vector (Mikolov et al., 2013b; Olah et al., 2020; Elhage et al., 2022b). However, these results are almost always for binary or categorical features, unlike the continuous features of space or time.
@@ 4.1 Verification via Generalization
Block holdout generalization
@@
We find that while generalization performance suffers, especially for the spatial datasets, it is clearly better than random. By plotting the predictions of the held out states or countries in Figures 11 and 12, a qualitatively clearer picture emerges. That is, the probe correctly generalizes by placing the points in the correct relative position (as measured by the angle between the true and predicted centroid) but not in their absolute position. We take this as weak evidence that the probes are extracting explicitly learned features by the model, but are memorizing the transformation from model coordinates to human coordinates. However, this does not fully rule out the underlying binary features hypothesis, as there could be a hierarchy of such features that do not follow country or decade boundaries.
@@
Cross entity generalization
@@
Similar to the above, we distinguish these hypotheses by training a series of probes where the train-test split is performed to hold out all points of a particular entity class. [^1] We only do this entities which do not make up the majority of the training data (e.g., as is the case with populated places for the world dataset and songs for the entertainment dataset) which is partially responsible for the discrepancies in the nominal cases for Tables 3 and 4. Table 4 reports the proximity error for the entities in the default test split compared to when heldout, averaged over all such splits as before. The results suggest that the probes largely generalize across entity types, with the main exception of the entertainment dataset. [^2] We note in this case the Spearman correlation is still high, suggesting this is an issue with bias generalization, as the different entity types are not uniformly distributed in time.
@@ 5 Space and Time Neurons
We also perform a series of neuron ablation and intervention experiments in Appendix B to verify the importance of these neurons in spatial and temporal modeling.
@@ 6 Related Work
Linguistic Spatial Models
Prior work has shown that natural language encodes geographic information (Louwerse & Zwaan, 2009; Louwerse & Benesh, 2012) and that relative coordinates can be approximately recovered with simple techniques like multidimensional scaling, co-occurrence statistics, or probing word embeddings (Louwerse & Zwaan, 2009; Mikolov et al., 2013a; Gupta et al., 2015; Konkol et al., 2017). However, these studies only consider a few hundred well known cities and obtain fairly weak correlations. Most similar to our work is (Liétard et al., 2021) who probe word embeddings and small language models for the coordinates of global cities and whether countries share a border, but conclude the amount of geographic information learned is “limited,” likely because the largest model they study was 345M parameters (500x smaller than Llama 70B).
@@
We consider a spatiotemporal model to be a necessary ingredient within a larger world model. The clearest evidence that such models are learnable from next-token prediction comes from GPT-style models trained on chess (Toshniwal et al., 2022) and Othello games (Li et al., 2022) which were shown to have explicit representations of the board and game state, with further work showing these representations are linear (Nanda et al., 2023). In true LLMs, Li et al. (2021) show that an entity’s dynamic properties or relations can be linearly read out from representations at different points in the context. Abdou et al. (2021) and Patel & Pavlick (2021) show LLMs have representations that reflect perceptual and conceptual structure in color and spatial domains.
@@
Interpretability
More broadly, our work draws upon many results and ideas from the interpretability literature (Räuker et al., 2023), especially in topics related to probing (Belinkov, 2022), BERTology (Rogers et al., 2021), the linearity hypothesis and superposition (Elhage et al., 2022b), and mechanistic interpretability (Olah et al., 2020). More specific results related to our work include Hanna et al. (2023) who find a circuit implementing greater-than in the context of years, and Goh et al. (2021) who find “region” neurons in multimodal models that resemble our space neurons.
@@ 7 Discussion
We have demonstrated that LLMs learn linear representations of space and time that are unified across entity types and fairly robust to prompting, and that there exists individual neurons that are highly sensitive to these features. We conjecture, but do not show, these basic primitives underlie a more comprehensive causal world model used for inference and prediction.
Our analysis raises many interesting questions for future work. While we showed that it is possible to linearly reconstruct a sample’s absolute position in space or time, and that some neurons use these probe directions, the true extent and structure of spatial and temporal representations remain unclear. We conjecture that the most canonical form of this structure is a discretized hierarchical mesh, where any sample is represented as a linear combination of its nearest basis points at each level of granularity. Moreover, the model can and does use this coordinate system to represent absolute position using the correct linear combination of basis directions in the same way a linear probe would. We expect that as models scale, this mesh is enhanced with more basis points, more scales of granularity (e.g. neighborhoods in cities), and more accurate mapping of entities to model coordinates (Michaud et al., 2023). This suggests future work on extracting representations in the model’s coordinate system rather than trying to reconstruct human interpretable coordinates, perhaps with sparse autoencoders (Cunningham et al., 2023).
We also barely scratched the surface of understanding how these spatial and temporal models are learned, recalled, and used internally, or to what extent these representations exist within a more comprehensive world model. By looking across training checkpoints, it may be possible to localize a point in training when a model organizes constituent is_in_place_X features into a coherent geometry or else conclude this process is gradual (Liu et al., 2021). We expect that the model components which construct these representations are similar or identical to those for factual recall (Meng et al., 2022a; Geva et al., 2023).
Finally, we note that the representation of space and time has received much more attention in biological neural networks than artificial ones (Buzsáki & Llinás, 2017; Schonhaut et al., 2023). Place and grid cells (O’Keefe & Dostrovsky, 1971; Hafting et al., 2005) in particular are among the most well-studied in the brain and may be a fruitful source of inspiration for future work on LLMs.
Acknowledgements
The authors would like to thank Sam Marks, Eric Michaud, Ziming Liu, Janice Yang, and especially Neel Nanda for their helpful discussions and feedback. W.G. was supported by Dimitris Bertsimas and an Open Philanthropy early career grant through the course of this work.
@@ Appendix A Datasets
Appendix B Neuron Ablations and Interventions
To better understand the role of space and time neurons in LLMs, we conduct several neuron ablation and intervention experiments.
Time Intervention
We study the effect of intervening on a single time neuron (L19.3610; correlation with art and entertainment release date of 0.77) within Llama-2-7b. Given a prompt of the form <media> by <creator> was written in 19, we pin the activation of the time neuron on all tokens and sweep over a range of pinned values, and track the predicted probability of the top five tokens. Results are depicted in Figure 7 and show that just adjusting the time neuron activation can change the next token prediction in all cases.
Figure 7: Prediction of decade of publication for a famous song, movie, and book when a time neuron (L19.3610) is pinned to a particular value, compared to 9 random neurons within the same layer (L19.[0-8]) of Llama-2-7b.
Neuron Ablations
We also study the effect of zero ablating neurons, and the contexts for which the loss increases the most. For a subset of Wikipedia which includes articles corresponding to world places and contemporary art and entertainment, we first run Llama-2-7b as normal and record the loss. Then, for two space neurons and two time neurons, we run the model with the neuron activation pinned to 0 (we always pin exactly one neuron to 0). For each neuron, we report the top 10 contexts in which the loss most increased for the next token prediction in Tables 5-8.
context true token loss increase
Bom Jesus has a rather dry tropical savanna climate (Köppen Aw 2.107
line tropical monsoon/humid subtropical climate (Köppen Am 2.035
of . The highest temperatures are reached at the end of the dry season in March 1.960
8.9 °C. In January, the average temperature is 1 8 1.930
a Tropical wet-and-dry climate (Köppen climate classification Aw 1.876
Goroka has a relatively cool tropical monsoon climate (Köppen Am 1.854
ie range from 26.4 °F in January to 7 0 1.835
wet summers and warm, very wet winters (Köppen climate classification Am 1.807
tropical wet and dry climate/semi-arid climate (Köppen Aw 1.783
ably mild, tropical-maritime climate, (Köppen climate classification Aw 1.762
Table 5: Contexts with the highest loss when ablating space neuron L20.7573 from Llama-2-7b.
context true token loss increase
mark is Mount Etna, one of the tallest active volcanoes in Europe 1.971
2,800 years, making it one of the oldest cities in Europe 1.676
keeper with 63 caps for Portugal including participation in the 198 4 1.631
15. Tenerife also has the largest number of endemic species in Europe 1.512
Georgia to the south-west. Mount Elbrus, the highest mountain in Europe 1.246
At one point, the village boasted the longest aluminium rolling mill in Western 1.219
centre and the leading economic hub of the Iberian Peninsula and of Southern 1.181
atican City, a sovereign state—and possibly the second largest in Europe 1.103
name. This is because the British Isles were likely repopulated from the I 1.082
Category 4 stadium by UEFA, hosted matches at the 199 8 1.072
Table 6: Contexts with the highest loss when ablating space neuron L20.7423 from Llama-2-7b.
context true token loss increase
was released in June 1993 as the fourth single from the album He 2.254
was released in November 1992 as the second single from her album He 1.973
93. Yearwood’s version was the third single from her album He 1.749
Hot Country Singles & Tracks chart, behind Shania Twain’s ¨ Any 1.574
was released in February 1992 as the third single from the album What 1.559
and provided additional production on her singles ”Like A Prayer” and ” Express 1.481
was released in May 1992 as the fourth single from the album What 1.367
filmmaker Ramesh Aravind in Telugu cinema. P. L 1.328
993 by record label Columbia as the second single from their second studio album Gold 1.272
Gessle for the duo’s 1991 album, Joy 1.253
Table 7: Contexts with the highest loss when ablating time neuron L18.9387 from Llama-2-7b.
context true token loss increase
016 as the third single from Reyes debut studio album, Louder ! 1.082
as the second radio single in support of the band’s third studio album, Life 0.996
Song, but ultimately lost the award to Barbra Streisand’s ” The 0.965
Released as the first single from the group’s seventh studio album, Super 0.961
original 30 Carry On films (1958–19 7 0.912
. After listening to No Doubt’s 2002 single ” Under 0.867
ix9ine for his debut mixtape Day69 (201 8 0.866
BA. It was originally featured on the group’s fifth studio album, The Album 0.864
ck that appeared in the Porky Pig cartoons It’s an 0.805
was written by Andrew Lloyd Webber and Tim Rice and produced by Fel 0.805
Table 8: Contexts with the highest loss when ablating time neuron L19.3610 from Llama-2-7b.
Appendix C Additional Results